OpenStudy (barrelracing):

help pretty please? Solve the following system of equations and show all work. y = 2x2 y = –3x −1

OpenStudy (barrelracing):

@Michele_Laino

OpenStudy (barrelracing):

i got this 2x² + 3x + 1 = 0 2x² + 2x + x + 1 = 0 2x(x+1)+1(x+1) = 0 (2x+1)(x+1) = 0 2x + 1 = 0 2x = -1 x = -1/2 y = 2(-1/2)² 2(1/4) 1/2 x + 1 = 0 x = -1 y = 2(-1)² 2(1) 2 (-1/2) (1/2) (-1, 2)

OpenStudy (michele_laino):

your quadratic equation is correct! nevertheless I got these solutions: \[\Large {x_1} = - \frac{1}{2},\quad {x_2} = - 1\]

OpenStudy (barrelracing):

ok thank you

OpenStudy (barrelracing):

can you help with another?